OpenStudy (anonymous):

Please help and im lost on these type of problems Determine the best method to solve the system of equations. Then solve the system. 1/4x+3/4y=1 2x+6y=8 A.substitution B. Elimination using multiplication; no solution C. Elimination using subtraction D. Elimination using multiplication; infinitely many solutions

OpenStudy (nikato):

Any ideas at all?

OpenStudy (anonymous):

im thinking c

OpenStudy (anonymous):

nope not really :(

OpenStudy (nikato):

Nice try but no.

OpenStudy (nikato):

By just looking at the problem, you can get rid of A and C. Because with substitution, it would be nice to have a variable with no coefficient in either one of the equations, which you don't have. And subtraction would be ugly because you'll be subtracting fractions. So now you're left with B or D with is elimination by multiplication

OpenStudy (anonymous):

might i make a suggestion?? \[1/4x+3/4y=1\\ 2x+6y=8 \] multiply the first equation by \(8\) and see what you get

OpenStudy (nikato):

Is the first equation 1/4x + 3/4y=1 Or (1/4)x + (3/4) y = 1

OpenStudy (anonymous):

1/4x+3/4y=1x8 right

OpenStudy (nikato):

No. What's (1/4x) (8)? And (3/4x) (8)?

OpenStudy (anonymous):

its 1/4x+3/4y=1

OpenStudy (anonymous):

i got 6x @nikato

OpenStudy (nikato):

Oops. Typo. If you multiply both sides by 8, you get have 8(1/4x + 3/4y) = 1(8) So (1/4x) 8 + (3/4y) 8 = 8 Can you simplify the left side?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

im starting to guess its B. or D.

OpenStudy (anonymous):

im guessing D.... am i right @nikato

OpenStudy (nikato):

That is absolutely correct.!

OpenStudy (nikato):

Because if you were to simplify that^, You would get 2x+6y=8 Which is exactly the same as the second equation. Since they are the SAME equation, they would share/have all the same answers